Use the table below to answer questions 8 through 10. Output (Q TP) TVC TC $80 100 $ 0 0 20 1 2 30 110 3 50 4 160 5 120 200 8 please answer 8-10. thank you!
Answer #1

8.

FC = TC when output is 0,

Here FC = 80

Here if firm does not operate, it will incur loss of 80

So option a is correct

9.

TC = TVC + FC

When 3 units are produced

TC = 50 + 80 = 130

So option d is correct

10

AVC = VC / Q

When 4 untis of output is produced

VC = TC - FC = 160 - 80 = 80

AVC = 80 / 4 = 20

So option b is correct